Michelle Payne deposited $7,800 in a savings account paying 7.00% simple interest. How long will it take for her investment to amount to $10,400?

Answer by ankor@dixie-net.com(22740)

Michelle Payne deposited $7,800 in a savings account paying 7.00% simple interest.
How long will it take for her investment to amount to $10,400?
:
simple interest formula: I = PRT
Find the amt of interest. 10400 - 7800 = 2600
7800*.07*t = 2600
546t = 2600
t = 2600/546
t = 4.76 ~ 5 years